# Learning to Play Two-Player Perfect-Information Games without Knowledge

### v4

@@ 7.3.2 Results
Table 7: Winning percentages against Mohex 3HNN of $\mathrm{UBFM}_{\mathrm{s}}$ using the learned evaluation function of Section 7.3 (the search time per action is the same for each player ; default settings are used for Mohex ; as many matches in first as in second player).
@@ 8 Conclusion
Firstly, we have generalized tree bootstrapping (tree learning) and shown that learning the values of the game tree instead of just learning the value of the root significantly improves learning performances with (iterative deepening) alpha-beta, and that this combination is more efficient than MCTS with tree learning, root learning or terminal learning, in the context of reinforcement learning without knowledge based on non-linear functions.
@@
Secondly, we have introduced the descent algorithm which explores in the manner of unbounded best-first minimax, intended to be used during the learning process. Unlike the latter, descent iteratively explores the sequences of best actions up to the terminal states. Its objective is to improve the quality of the data used during the learning phases, while keeping the advantages of unbounded best-first minimax. In the context of our experiments, the use of descent gives significantly better performances than the use of alpha-beta, unbound minimax, or MCTS.
@@
Thirdly, we have suggested to replace the classic gain of a game (+1/1+1/-1+ 1 / - 1) by a different terminal evaluation function. We have propose different general terminal evaluations, such as the depth heuristic, which takes into account the duration of games in order to favor quick wins and slow defeats. Our experiments have shown that the use of a reinforcement heuristic improves performances and that the depth heuristic is a useful general heuristic.
@@
We have also proposed a new action selection distribution which does not take into account the value of states but only their order.
In the context of this experiments, the techniques that we propose are the best approach of reinforcement learning of state values. We apply all these different techniques to design program-players to Hex with swap (size 11 and 13) surpassing, for the first time, the level of Mohex 3HNN (the best Hex program based on supervised learning and dedicated Hex algorithms and knowledge) by reinforcement learning from self-play without knowledge (or Monte Carlo techniques).
Overall, the results show that unbounded minimax, which has obtained good results in the experiments of this article, has been under-studied in the literature and that unbounded minimax and its variants, that we offer, seems to constitute particularly efficient alternatives to alpha-beta and MCTS, in the era of reinforcement learning.
A promising research perspective is the parallelization of $\mathrm{UBFM}$ and descent when learning: $\mathrm{UBFM}$ determines the next action to play and descent determines the pairs to learn. The other research perspectives include the application of our contributions to the game of Go and to General Game Playing (at first with perfect information). They also include the application and adaptation of our contributions to the contexts of hidden information, stochastic and multiplayer games. Moreover, they include the application of our contributions to optimization problems, such that the RNA Inverse Folding problem [63].
@@
The realized programs (based on the descent framework) are programmed in Python. Switching to a faster performing language should reduce the learning time by a factor between two and five.
Note to conclude that, on the one hand, the descent framework with the unbounded minimax faced Polygames during the computer olympiad of 2020 and beat it at Othello 8x8, Othello 10x10 and Breakthrough. In fact, 5 gold medals were won by the proposed algorithms for the following games: Othello 10x10, Breakthrough, Surakarta, Amazons, and Clobber. At Clobber and at Surakarta, one of the participating programs was also based on AlphaZero. That of Clobber used in addition the combinatorial game theory to solve states. Moreover, a silver medal were won by the proposed algorithms at Othello 8x8. It is the first time at the Computer Olympiad that the same program wins 5 gold medals.
On the other hand, the descent framework has also competed in the 2021 Computer Olympiad. This time it won $11$ gold medals (Hex $11\times 1$ 11, Hex $13\times 13$, Hex $19\times 19$, Havannah $8\times 8$, Havannah $10\times 10$, Othello $8\times 8$, Surakarta, Amazons, Breakthrough, Brazilian Draughts, Canadian Draughts; there was no competition at Othello 10x10 and Clobber). Descent notably beat Polygames at games where they met (Hex $13\times 13$, Hex $19\times 19$, Havannah $8\times 8$, Havannah $10\times 10$).

### v5

@@ 9.4.2 Results
Table 8: Winning percentages against Mohex 3HNN of $\mathrm{UBFM}_{\mathrm{s}}$ using the learned evaluation function of Section 9.4 (the search time per action is the same for each player ; default settings are used for Mohex ; there are as many matches in first as in second player).
10 Application at Othello
In this section, we apply Athénan on the game of Othello. After completing the reinforcement learning process without knowledge, we evaluated Athénan against the state-of-the-art at Othello: the Edax program [^10] Edax version 4.4: https://github.com/abulmo/edax-reversi. We start by presenting the state-of-the-art at Othello (Section 10.1), then we describe the experiment carried out (Section 10.2) and its technical details (Section 10.3), and finally we present the results (Section 10.4).
10.1 Othello Related Work
Although Othello is one of the first games in which artificial intelligence has reached a superhuman level, there is still research on Othello. Edax and Saio are the Othello state-of-the-art programs [67, 68, 69, 70]. Saio is not free and is similar to Edax. Edax is an open source highly optimized program. It is based on minimax and dedicated Othello techniques: multi-probcut tree search, tabular value functions (as evaluation function) generated from expert knowledge, and opening sequences [71].
There are other programs at Othello.
A move predictor [69] was designed at Othello using a convolutional neural network learned by supervised learning. However, its level is low. In particular, it only beats Edax when Edax plans at depth 2 (which takes it on the order of $10^{-4}$ seconds), whereas Edax plans with a depth between 20 and 30 for a search time of the order of seconds.
AlphaZero has been applied recently at Othello [68]. This AlphaZero program, named Olivaw, is as strong as Edax with a search depth of 8. Their program reached a high level at Othello, and was able to beat a national champion at Othello.
See [67, 68, 69] for a detailed Related Work about Computer Othello.
10.2 Experiment
We now apply Athénan to carry out a long self-play reinforcement learning at Othello. More precisely, we use completed Descent (Algorithm 10) with tree learning (Algorithm 3), completed ordinal distribution (see Section 5 and Algorithm 14), the game score as reinforcement heuristic (see Section 6), and the stratified experience replay (Section 8).
10.3 Technical Details
[…]
Search time per action is $\tau=5s$. The stratified experience replay parameters used are: the batch size $B=4000$, the memory size $\mu=100$, the duplication factor $\delta=3$
We use the following neural network as adaptative evaluation function: a residual network with a convolutional layer with $373$ filters, followed by $8$ residual blocks (two $3\times 3$ convolutions per block with $373$ filters each following with a squeeze-and-excitation layer [72] whose ratio is $16$), followed by a $1\times 1$ convolution with $4453$ filters, followed by a global sum pooling [73], followed by a flat layer, and followed by a fully connected hidden layers (with $4453$ neurons). The activation function used is the ReLU. Number of weights of the neural network is $41,695,147$. he Adam parameters are $\lambda=0.0001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-5}$. The parameter of the $L_{2}$ regularization is $0.99$.
Moreover, when the children of a state are evaluated by the neural network, they are batched and thus evaluated in parallel (on the only GPU). The evaluation function has been pre-initialized by learning the values of random terminal states (of the order of $10,000,000$). Resolved states are kept in memory (the memory of the resolved states is emptied every 12 learning hours). TNo simulated annealing was used for the ordinal distribution. The parameter of the ordinal distribution is a random number drawn uniformly from [0,1] each time this distribution is used. The other settings are the same as those in Section 3.3.1.
The reinforcement learning process lasted $52,470$ matches. Note that the number of data used during the learning process is of the order of $426\cdot 10^{7}$, the number of neural network evaluations is of the order of $606\cdot 10^{6}$, and the number of state evaluations is of the order of $3.49\cdot 10^{9}$.
10.4 Results
The winning percentages against Edax of Athénan (i.e. of $\mathrm{UBFM}_{\mathrm{s}}$ using the learned evaluation function generated by Descent) are described in Table 9. Thus, the proposed techniques have made it possible to exceed the level of Edax at Othello. Recall that this reinforcement learning was carried out without using prior knowledge about Othello strategies.
search time win - loss win draw loss $95\%$ confidence radius total matchs
$1.5s$ $20\%$ $39.37\%$ $41.25\%$ $19.37\%$ $3,4\%$ $800$
$15s$ $12.25\%$ $25.37\%$ $61.5\%$ $13.12\%$ $3,4\%$ $800$
Table 9: Winning percentages against Edax of $\mathrm{UBFM}_{\mathrm{s}}$ using the learned evaluation function of Section 10.2 (the search time per action is the same for each player ; there are as many matches in first as in second player).
11 Application at Arimaa
In this section, we apply Athénan on the game of Arimaa and we evaluate the performed process of reinforcement learning without knowledge against the state-of-the-art of Arimaa: the Sharp program [74]. We start by presenting the state-of-the-art at Arimaa (Section 11.1), then we describe the experiment carried out (Section 11.2) and its technical details (Section 11.4), and finally we present the results (Section 11.5).
11.1 Arimaa Related Work
Since 2015, state-of-the-art at Arimaa is the program Sharp [74]. Sharp combined traditional alpha–beta pruning with handcrafted heuristic functions. But it uses a lot of improvements of alpha-beta, called killer moves, history heuristic, quiescence search, extensions, late move reduction, (theses techniques required dedicated knowledge to be used). It also uses techniques dedicated to Arimaa: specific algorithms and knowledge used by alpha-beta and its improvements (mainly for move ordering, pruning, and state evaluating).
Alternative approaches to alpha-beta at Arimaa have been studied, notably based on Monte Carlo, but without success [74]. There have also been some attempts to apply AlphaZero at Arimaa, but they have all failed to our knowledge [75].
11.2 Experiment
We now apply Athénan to carry out a long self-play reinforcement learning at Arimaa. More precisely, we use completed Descent (Algorithm 10) with tree learning (Algorithm 3), completed ordinal distribution (see Section 5 and Algorithm 14), the stratified experience replay (Section 8), and the Arimaa reinforcement heuristic described in the following section.
11.3 Arimaa Reinforcement Heuristic
The reinforcement heuristic (see Section 6) we used for Arimaa is a variation of the presence heuristic (Section 6) which takes into account the importance of the pieces which comes directly from the rules.
We use the following state evaluation function, denotedy by $\mathrm{h}_{\mathrm{arimaa}}$, as reinforcement heuristic. Let $s$ be a game state. Then, $\mathrm{h}_{\mathrm{arimaa}}\left(s\right)=\begin{cases}69+V\left(s\right)&\text{if }\text{first player wins}\\ -69+V\left(s\right)&\text{if }\text{second player wins}\end{cases}$ where $P_{1}\left(s\right)$ (resp. $P_{2}\left(s\right)$) the set of pieces of the first player (resp. second player) in $s$, $V\left(s\right)$ is the difference of the values of the two players pieces:
$V\left(s\right)=\sum_{p\in P_{1}\left(s\right)}v_{\mathrm{arimaa}}\left(p\right)-\sum_{p\in P_{2}\left(s\right)}v_{\mathrm{arimaa}}\left(p\right)\text{,}$
and $v_{\mathrm{arimaa}}\left(p\right)$ is the value the piece $p$ corresponding to the importance of that piece according to the rule. The value $v_{\mathrm{arimaa}}\left(p\right)$ is defined as follows: $v_{\mathrm{arimaa}}\left(\mathrm{elephant}\right)=5$, $v_{\mathrm{arimaa}}\left(\mathrm{camel}\right)=4$, $v_{\mathrm{arimaa}}\left(\mathrm{horse}\right)=3$, $v_{\mathrm{arimaa}}\left(\mathrm{dog}\right)=2$, $v_{\mathrm{arimaa}}\left(\mathrm{cat}\right)=1$, $v_{\mathrm{arimaa}}\left(\mathrm{rabbit}\right)=6$.
11.4 Technical Details
[…]
Search time per action is $\tau=5s$. The stratified experience replay parameters used are: the batch size $B=4000$, the memory size $\mu=100$, the duplication factor $\delta=3$
We use the following neural network as adaptative evaluation function: a residual network with a convolutional layer with $263$ filters, followed by $4$ residual blocks (two $3\times 3$ convolutions per block with $263$ filters each), followed by a $1\times 1$ convolution with $2217$ filters, followed by a global sum pooling [73], followed by a flat layer, and followed by a fully connected hidden layers (with $2217$ neurons). The activation function used is the ReLU. Number of weights of the neural network is $10,511,017$. The Adam parameters are $\lambda=0.0001$, $\beta_{1}=0.9$, $\beta_{2}=0.99$ and $\epsilon=10^{-7}$.
Moreover, when the children of a state are evaluated by the neural network, they are batched and thus evaluated in parallel (on the only GPU). The evaluation function has been pre-initialized by learning the values of random terminal states (of the order of $10,000,000$). Resolved states are kept in memory (the memory of the resolved states is emptied every 12 learning hours). The other settings are the same as those in Section 3.3.1.
The reinforcement learning process lasted $6,317$ matches. Note that the number of data used during the learning process is of the order of $768\cdot 10^{7}$, the number of neural network evaluations is of the order of $10^{9}$, and the number of state evaluations is of the order of $30\cdot 10^{9}$.
11.5 Results
The winning percentages against Sharp of Athénan (i.e. of $\mathrm{UBFM}_{\mathrm{s}}$ using the learned evaluation function generated by Descent) are described in Table 10. Thus, the proposed techniques have made it possible to exceed the level of Sharp at Arimaa. Recall that this reinforcement learning was carried out without using prior knowledge about Arimaa strategies.
search time win loss $95\%$ confidence interval total matchs
$2.5s$ $90.33\%$ $9.66\%$ $4\%$ $600$
$10s$ $92.5\%$ $7.5\%$ $4\%$ $600$
Table 10: Winning percentages against Sharp of $\mathrm{UBFM}_{\mathrm{s}}$ using the learned evaluation function of Section 10.2 (the search time per action is the same for each player ; there are as many matches in first as in second player).
12 Application at Morpion Solitaire
In this section, we apply Athénan on the game of Morpion Solitaire (also called Join Five), a puzzle game (one-player game). We start by detailing the rules of Morpion Solitaire (Section 12.1), then presenting the state-of-the-art at Morpion Solitaire (Section 12.2). Next, we describe the experiment carried out (Section 12.3) and its technical details (Section 12.4), and finally we present our results (Section 12.5).
12.1 Morpion Solitaire Rules
The goal of the Morpion Solitaire is to play as long as possible. The score of the Morpion Solitaire is therefore the number of moves played since the start of the game. The Morpion Solitaire board is an infinite size square board. At the start of the game, some points are already placed in the shape of a cross. At each turn, the player places a point on the board and crosses out an alignment of 5 points not already crossed out in that direction. A point can therefore only be crossed out at most 4 times (horizontally, vertically, diagonally and antidiagonally). The 5 points must include the point just placed. A point can only be placed if it will belong to an alignment of 5 points that can be crossed out. A move includes the point placement plus the choice of the 5 aligned points and their crossing-out. As soon as the player cannot place any more points, the game is over. The initial board and the board of an example of first move are represented in Figure 17.
Figure 17: Initial board for Morpion Solitaire (left) ; The board of Morpion Solitaire after a first move (right).
The Morpion Solitaire has a variant, called the touching version (the Morpion Solitaire, described above, is called the disjoint version). With the touching variant, the endpoints of an already drawn line can be used as an endpoint for a new line aligned in the same direction. This variant is not studied in this paper.
12.2 Morpion Solitaire Related Work
An upper bound on the maximum score at Morpion Solitaire (disjoint version) is $84$ [76]. The human record is 68 [77]. The best actual record is 82. The record of 82 was obtained with Nested Rollout Policy Adaptation [78] and Beam Nested Rollout Policy Adaptation [79], which are Monte Carlo algorithms, using online learning, dedicated to one-player games.
AlphaZero has been applied with the Ranked Reward technique to Morpion Solitaire [77]. The record obtained by this AlphaZero program is only $67$.
ExIt, the reinforcement learning algorithm without knowledge, has been combined with Nested Rollout Policy Adaptation and applied to Morpion Solitaire [80]. The record of this ExIt program is $73$.
The touching version has also been studied. The record for this variant of Morpion Solitaire is 178 [81].
12.3 Experiment
We now apply Athénan to carry out a long self-play reinforcement learning without knowledge at Morpion Solitaire. More precisely, we use Descent (Algorithm 9) with tree learning (Algorithm 3), ordinal distribution (see Section 7), the stratified experience replay (Section 8), and the score heuristic as reinforcement heuristic (Section 6).
Remark 20.
Note that completion is not used since it only concerns two-player games. However, resolved states are still used. In this context, a state is resolved only if all its children are resolved.
12.4 Technical Details
All endgame are draw unless the score of 84 is reached (win).
[…]
Search time per action is $\tau=1s$. The stratified experience replay parameters used are: the batch size $B=3000$, the memory size $\mu=100$, the duplication factor $\delta=3$.
We use the following neural network as adaptative evaluation function: a residual network with a convolutional layer with $186$ filters, followed by $8$ residual blocks (two $3\times 3$ convolutions per block with $186$ filters each, following with a squeeze-and-excitation layer [72] whose ratio is $16$), followed by a flat layer, and followed by two fully connected hidden layers (with each $104$ neurons). The activation function used is the ReLU. Number of weights of the neural network is $9,989,285$. The Adam parameters are $\lambda=0.0001$, $\beta_{1}=0.9$, $\beta_{2}=0.99$ and $\epsilon=10^{-8}$.
Moreover, when the children of a state are evaluated by the neural network, they are batched and thus evaluated in parallel (on the only GPU). The evaluation function has been pre-initialized by learning the values of random terminal states (of the order of $10,000,000$). Resolved states are kept in memory (the memory of the resolved states is emptied every 12 learning hours). The other settings are the same as those in Section 3.3.1.
The reinforcement learning process lasted $58,971$ matches. Note that the number of data used during the learning process is of the order of $126\cdot 10^{7}$, the number of neural network evaluations is of the order of $0.19\cdot 10^{9}$, and the number of state evaluations is of the order of $1.8\cdot 10^{9}$.
12.5 Results
The evolution of the score obtained in Morpion Solitaire using Athénan (i.e. using $\mathrm{UBFM}_{\mathrm{s}}$ based on the evaluation function generated by Descent) with $0.1$ second of search per action is described in Figure 18. The score of $82$ is reached in the $58,971$ match.
Thus, the state-of-the-art level has been reached with a more general algorithm than the state-of-the-art techniques on this game. Recall that this result was not achieved with ExIt and AlphaZero, the two alternative algorithms of Athénan.
This result further shows that Athénan is applicable to single-player games.
Figure 18: Evolution of the score obtained in Morpion Solitaire using the evaluation function learned and $\mathrm{UBFM}_{\mathrm{s}}$ with $0.1$ second of search per action.
13 Computer Olympiad Results
In this section, we briefly present Athénan’s results in the global board game artificial intelligence competition, namely the Computer Olympiad.
Athénan won a lot of gold medals at the Computer Olympiad: 48 in total. It won eleven gold medals in 2024 [82], sixteen gold medals in 2023 [83], five gold medals in 2022, eleven in 2021, and again five in 2020 [84].
Moreover, Athénan is currently the defending champion on 17 games (the eleven 2024 gold medals plus 6 previous uncontested gold medals). All Athénan results are summarized in Table 11.
Note that, no other program has exceeded 5 gold medals in the same year since the beginning of this competition in 1989. In fact, reaching 5 medals so far was an exceptional achievement.
2020 2021 2022 2023 2024
Amazons Gold Gold 1 Gold Gold
Breakthrough Gold Gold Gold Gold Gold
Clobber Gold 1 1 Gold Gold
Surakarta Gold Gold Gold Gold Gold
Othello 8x8 Silver Gold Silver Bronze
Othello 10x10 Gold 1 1 Gold 1
Othello 16x16 Gold
Hex 11x11 Gold 1 Gold Gold
Hex 13x13 0 Gold 1 Gold Gold
Hex 19x19 0 Gold 1 Gold 1
Havannah 8 0 Gold 1 Gold Gold
Havannah 10 Gold 1 Gold 1
Canadian Draughts Gold Gold Gold 1
Brazilian Draughts Gold 0 Silver
International Draughts Silver 0 Silver
Connect6 Silver Silver Silver 0
Outer-Open-Gomoku Bronze Silver Bronze 0
Ataxx Gold Gold Gold
Santorini Gold Gold 0
Lines of Action 0 Gold Gold
Xiangqi 0 Gold 1
Arimaa Gold 1
Shobu Gold
Table 11: Athénan Computer Olympiad results (00: participation without result ; $1$: participation without opponent, i.e. title not contested ; empty: no participation).
@@ 14.1 New algorithms: Athénan components
Firstly, we have generalized tree bootstrapping (tree learning) in the context of reinforcement learning without knowledge based on non-linear functions. We have shown that learning the values of the game tree instead of just learning the value of the root significantly improves learning performances.
@@
Secondly, we have introduced the Descent algorithm which explores in the manner of Unbounded Minimax, intended to be used during the learning process. Unlike Unbounded Minimax, Descent iteratively explores the sequences of best actions up to terminal states. Its objective is to improve the quality of the data used during the learning phases, while keeping the advantages of Unbounded Minimax. In the context of our experiments, the use of Descent gives strongly better performances than the use of Alpha-Beta, Unbounded Minimax, or MCTS.
Thirdly, we proposed the completion technique which allows to take into account the resolution of states, notably in the context of Unbounded Minimax and Descent. Note that the impact assessment of this technique was carried out in [35]. This technique improves performance but the gain is lower than the techniques evaluated in this article.
@@
Fourthly, we have suggested to replace the classic gain of a game (+1/1+1/-1+ 1 / - 1) by different terminal evaluation functions, called reinforcement heuristics. We have propose different general terminal evaluations, such as the depth heuristic, which takes into account the duration of games in order to favor quick wins and slow defeats. Our experiments have shown that the use of a reinforcement heuristic improves performances. Our study recommends using the score heuristic when the game has one and otherwise using the depth heuristic.
@@
Fifth, we have proposed a new action selection distribution which does not take into account the value of states but only their order.
We then combined all these techniques within Athénan, a zero-knowledge reinforcement learning algorithm, like AlphaZero or ExIt.
14.2 Evaluation of Athénan
In the unpublished initial version of this article [37], Athénan has not been compared to the two state-of-the-art zero-knowledge reinforcement learning algorithms: AlphaZero and ExIt. In the meantime, Athénan’s comparison with AlphaZero was realized in collaboration with Tristan Cazenave in another article [22]. In this external study, Athénan with the classic game gain as reinforcement heuristic is at least 7 times faster than AlphaZero on equal hardware. It is also at least 30 times faster using the reinforcement heuristics presented in this article. Athénan with one GPU even outperforms AlphaZero with 100 GPUs on some games.
In the new version of this paper, we have therefore added the missing experience: the comparison with ExIt. In the context of the experiments conducted in this article, Athénan performs much better than ExIt. ExIt’s average performance does not even reach in 15 days the performance that Athénan obtained in one day. Note that in this study, Athénan only used the classic game gain as reinforcement heuristic. Thus, we should achieve a result 4 times higher with the advanced reinforcement heuristics proposed in this article.
We also showed that Athénan arrived, without using any knowledge [^11] Without using any knowledge other than the rules of the game or arising trivially from the rules of the game (namely the use of board symmetry, encoding of board edge membership in Hex, order of importance of pieces for the Arimaa reinforcement heuristic)., to surpass the state-of-the-art of many games, namely Hex size 11 and size 13, Othello, and Arimaa. This contrasts with the outperformed state-of-the-art programs which use dedicated knowledge.
Moreover, Athénan achieved the state-of-the-art level at Morpion Solitaire. Note that the state-of-the-art record at Morpion Solitaire was achieved using algorithms restricted to single-player games. Conversely, this result shows that in addition to being applicable to two-player games, Athénan is also applicable to single-player games. This is all the more noteworthy since attempts to achieve state-of-the-art level at Morpion Solitaire with ExIt and AlphaZero failed.
Finally, we briefly presented Athénan’s results at the Computer Olympiad, the global artificial intelligence board game competition. Athénan broke all records at this competition, tripling the record of gold medals obtained in a single year (reaching 16 gold medals), winning a total of 48 gold medals in 5 years since its first participation, and being the defending champion on 17 games.
Overall, the results show that Athénan and its components constitute particularly efficient and above all effective alternatives to other approaches based on Alpha-Beta and MCTS, in the era of reinforcement learning.
@@ 14.2 Evaluation of Athénan
Athénan has been programmed in Python. Switching to a faster performing language should reduce the learning time by a factor between two and five and will increase the winning percentages during confrontations.
14.3 Perspectives
We did not evaluate the ordinal distribution in this article. We will see in a future article that this techniques improves average performance. However, its significant gain is much lower than the other techniques proposed in this article (except completion).
Work is underway on the generalization of these algorithms in the following contexts: stochastic games [85], multiplayer games [51], and usage of a policy.
A promising research perspective is the parallelization of Unbounded Minimax and Descent when learning: Unbounded Minimax determines the next action to play and Descent determines the pairs to learn (by batching evaluations on a single GPU or using two GPUs).
Another promising perspective is to perform training using Athénan with two neural networks playing against each other. This could potentially reduce numerical problems and the risks of local optimums. To avoid losing learning speed, on the one hand, each player performs a Descent search during its opponent’s turn. On the other hand, the neural networks are evaluated in parallel, on the same graphics card (or on two different ones). This approach can also be generalized to a pool of neural networks performing matches in parallel. In this context, it might be more interesting to replace the ordinal distribution with the max distribution. These approaches should speed up training and increase efficiency, but they will reduce effectiveness.
The other research perspectives include the application of our contributions to General Game Playing (at first with perfect information). A yet-to-be-determined Athénan parameterization, perhaps with the use of small neural networks, could provide interesting performance in the context of General Game Playing. A modification of Athénan that could be interesting in this context is to perform "online" learning, i.e. instead of performing the learning phase after each match, to perform it after each search, or even after each Descent iteration, or maybe even after each updating of game tree nodes.
The other research perspectives also include the application and adaptation of our contributions to the contexts of hidden information.
Finally, they include the application of our contributions to optimization problems, such that the RNA Inverse Folding problem [86].